OpenStudy (anonymous):

Can someone please help me? I have been trying to solve this for hours :( A triangular prism has an equilateral triangle as its base with sides that measure 2 yards and a height of 5 yards. What is the surface area of this prism if the height of the prism is 8 yards? 45 yd² 48 yd² 58 yd² 80 yd²

OpenStudy (anonymous):

try drawing a quick sketch, it might help you

OpenStudy (lgbasallote):

is it lateral? or total surface areaa?

OpenStudy (anonymous):

It's total surface area.

OpenStudy (anonymous):

That's the thing, I've tried doing things like drawing it, I just can't seem to get the hang of it. :(

OpenStudy (lgbasallote):

total surface area = lateral surface area + base area + base area lateral surface area = height x perimeter base

OpenStudy (lgbasallote):

does that help??

OpenStudy (phi):

Are you sure about the question? A triangular prism has an equilateral triangle as its base with sides that measure 2 yards and a height of 5 yards. You cannot have an equilateral triangle with sides = 2 and an altitude = 5 You can have an isosceles triangle with these measures. If we use 2 as the base and 5 as the height, the area of the bottom is 1/2 5*2= 5 sq yd the top is the same, so 10 total for top and bottom the 3 sides are each 2*8= 16 sq yd = 48 sq yd total 58 sq yd

OpenStudy (anonymous):

@lgbasallote YES! Thank you so much! Could you stick around for a second to check over my equation?

OpenStudy (phi):

oops, the 3 legs cannot be the same the 3 sides are each 2*8= 16 sq yd = 48 sq yd

OpenStudy (anonymous):

@phi Oh excellent! Thank you so much, you would not believe how much that helped. Thank you so much for explaining it and everything. :D

OpenStudy (phi):

I don't think I explained it. There is no way to get any of your answers that I can see.

OpenStudy (lgbasallote):

ok..go on

OpenStudy (phi):

If it is equilateral with sides=2, the area of a triangle is sqrt(3)

OpenStudy (phi):

if the triangle is a base 2 with height 5, we do not know the length of the other 2 sides. If we assume isosceles, the sides would be sqrt(26). Not good.